EXCITED STATE PROPERTIES OF PADDLEWHEEL COMPLEXES INVOLVING METAL-METAL (M = MO, W) QUADRUPLE BONDS SUPPORTED BY AMIDINATE AND CARBOXYLATE LIGANDS

\underline{BRIAN G. ALBERDING} , MALCOLM H. CHISHOLM, and TERRY L. GUSTAFSON , The Ohio State University, Department of Chemistry, Columbus, Ohio, 43210.

Spectroscopic studies of the trans-substituted compounds M_{2}(O_{2}CCH_{3})_{2}[(N^{i}Pr)_{2}CCCR]_{2} (where M = Mo or W and R = aromatic hydrocarbon) have been carried out. These complexes have been shown to possess optically active excited states in both the singlet and triplet manifolds that can be classified as either charge transfer or metal centered in character. The nature of the excited states has been assigned based on steady state absorbance and emission studies and supported by electronic structure calculations. Furthermore, excited state dynamics have been determined from both nsand fs-transient absorption spectroscopies. Comparisons will be made between molybdenum and tungsten.